Hi. I was wondering if somebody could help me with finding the absolute value of:
absolute value of 4-3x=19
Also, I need to find the solution of this inequality and graph on a number line:
-2 less than or equal to x-4 over 3 less than or equal to 4
I appreciate your time and help. Thank you.

Answer by lwsshak3(11628)

Hi. I was wondering if somebody could help me with finding the absolute value of:
absolute value of 4-3x=19
Also, I need to find the solution of this inequality and graph on a number line:
-2 less than or equal to x-4 over 3 less than or equal to 4
**
|4-3x|=19
two conditions would satisfy this equation:
(4-3x)=19
-3x=19-4=15
-x=15/3=5
x=-5
or
-(4-3x)=19
-4+3x=19
3x=19+4=23
x=23/3
ans:
x=-5
or
x=23/3
..
-2≤x-4/3≤4
multiply by 3
-6≤x-4≤12
add 4
-2≤x≤16
Ans:
Interval notation:
[-2,16]
number line:
<.......[.............]..........>
.......-2.....0....16